One train travels 50 miles per hour faster than a second train. If the slower train travels 225 miles in the same amount of time as the faster train travels 375 miles, what is the speed of the faster train?
Answer by mananth(16946)

Train A x mph
Train B 50 faster

Train A 225 miles
Train B 375 miles
Train A time =Train B time

225/x=375/(x+50)cross multiply
225(x+50)=375x
225x+11250=375 x
375x-225x=11250
150x=11250
/150
x=75 mph Train A

125 mph Train B